OpenStudy (anonymous):

find the derivative of f(x)= 12cosxsinx

OpenStudy (kainui):

Use the product rule here.

OpenStudy (anonymous):

12 cos 2x

OpenStudy (anonymous):

would the answe be 12cos^2xsin^2x?

OpenStudy (raden):

first, better convert 12cosxsinx to 6sin2x next find its derivative

OpenStudy (kainui):

Almost gnarballs, the answer is really close to that though, don't listen to these guys, they're giving poor advice.

OpenStudy (anonymous):

well since the d of cos is -sin would the whole thing be negative?

OpenStudy (kainui):

So the derivative of cosxsinx is 12[(-sinx)(sinx)+(cosx)(cosx)] and I left it this way to show you how the product rule is used.

OpenStudy (kainui):

So you can see it simplifies to 12cos^2x-12sin^2x, yes? You were nearly correct.

OpenStudy (anonymous):

oh the twelve distributes to both parts ! of course.... the program did not like that answer though

OpenStudy (anonymous):

\[\large{ f(x)= 12 \cdot \cos x \cdot\sin x}\] \[\large{ f'(x)= {d\over dx}(12 \cdot \cos x \cdot \sin x) \\ f'(x)= 12\cdot {d\over dx}(\cos x \cdot \sin x) \\ f'(x)= 12\cdot \left( (\cos x)'(\sin x) + (\cos x)(\sin x)' \right) }\]

OpenStudy (kainui):

It appears as though it simplifies with a trigonometric identity further http://www.wolframalpha.com/input/?i=d%2Fdx%2012cosxsinx&t=crmtb01

OpenStudy (anonymous):

no luck

OpenStudy (anonymous):

Oh right, this identity. http://www.youtube.com/watch?v=IaZFDmVtyyU

OpenStudy (kainui):

Perhaps the program you're typing into doesn't like the notation of squaring sine and cosine like that. Try: 12(cosx)^2-12(sinx)^2

OpenStudy (anonymous):

wow its my fault i changed it to x to post here, so i used x to answer it, while the constant they use is q ! thank you very much guys

OpenStudy (kainui):

lol Thankfully no trig identities needed, although they make it look nicer they tend to over complicate things sometimes.